# Scaling limit of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant boundary conditions

### v2

@@ 4.1 Continuous and discrete spectrum
A key result of this paper is the conjecture, above, which was motivated by our numerical work. It describes a certain one to one relation between $b(L)$ and $b_{*}(L)$. The former is computed via the Bethe roots corresponding to a state $|\Psi_{L}\rangle$ in a multiplet ${\cal M}_{\cal S}$, labeled by $2p=2{\cal S}+1-(n+2)$, the non-negative integer ${\tt d}$ and one of the $N({\tt d}\,|\,{\cal S})={\rm par}_{2}({\tt d})-{\rm par}_{2}({\tt d}-2{\cal S}-1)$ solution sets $\bm{w}=\{w_{a}\}_{a=1}^{\tt d}$ of the algebraic system (3.2). The notation $b_{*}(L)$ stands for a solution of the quantization condition (3.40) treated as an equation for $b(L)$ with the correction terms ignored. Accepting the conjecture to be true, one can determine the spectrum of $b(L)$ for the low energy states at large $L$ via a study of the quantization condition. In turn, the results allow one to characterize the spectrum of conformal dimensions as well as the space of low energy states in the scaling limit.
Consider the quantization condition (3.40) and suppose that the phase shift $\delta(\bm{w},p,s)|_{s\mapsto b(L)}\ll\log(L)$. Then the first term dominates and one can develop an asymptotic expansion for $b(L)$ in $1/\log(L)$. The leading and subleading asymptotic behaviour reads as
@@
Each low energy Bethe state in ${\cal H}_{L|{\cal S},S^{z}}^{({\rm cont})}$ is uniquely specified by the non-negative integer ${\tt d}$, a solution set $\bm{w}=\{w_{a}\}_{a=1}^{\tt d}$ of the algebraic system (3.2) with $s\mapsto b(L)$, as well as the even or odd integer $\mathfrak{m}$ that enters into the asymptotics (4.1). For $L\gg 1$ the value of $b_{\mathfrak{m}}(L)$ becomes densely distributed in the segment $(-b_{\mathfrak{m}_{\rm max}}(L),+b_{\mathfrak{m}_{\rm max}}(L))$. The density of states is obtained from the quantization condition (3.40) written in logarithmic form:
$\big{[}4s\log(L/L_{0})+\delta(\bm{w},p,s)\big{]}|_{s\mapsto b_{\mathfrak{m}}(L)}=2\pi\mathfrak{m}+O(L^{-\epsilon})\ .$ (4.4)
@@
Here the branch of the logarithm needed to define $\delta$ from the relation (3.41) is taken such that the phase shift is a continuous function of $s$ in the strip $|\Im m(s)|<\varepsilon$ for some $\varepsilon>0$ (it is being assumed that $\mbox{e}^{\frac{{\rm i}}{2}\delta}$ contains no zeroes or poles for real $s$). The term in the square brackets in the left hand side of (4.4) is a monotonic function of $s$ for $L$ sufficiently large. This way one concludes that the number of states in ${\cal H}_{L|{\cal S},S^{z}}^{({\rm cont})}$ with fixed ${\tt d}$ such that $\Re e\big{(}b(L)\big{)}$ lies in the interval $(s,s+\Delta s)\in(-b_{\mathfrak{m}_{\rm max}}(L),+b_{\mathfrak{m}_{\rm max}}(L))$ is given by $\rho^{({\tt d})}_{p}(s)\,\Delta s$ with [^6] This line of arguments is analogous to the standard derivation in the root density approach. One introduces a monotonic increasing counting function which evaluates to (half-)integers at the Bethe roots similar as the l.h.s of (4.4) evalutes to odd/even integers $\mathfrak{m}$ multiplied by $\pi$ when $s$ is swapped for $b_{\mathfrak{m}}(L)$ and $L\gg 1$. Differentiating the counting function in the root density approach yields the root density, while we obtain the density of states (4.5) by differentiating (4.4) and dividing by $4\pi$.
@@
$\rho^{({\tt d})}_{p}(s)=\frac{1}{\pi}\,N({\tt d}\,|\,{\cal S})\,\log\big{(}2^{\frac{n+2}{n}}\,L/L_{0}\big{)}+\frac{1}{2\pi{\rm i}}\ \partial_{s}\log\Bigg{[}\bigg{(}\frac{\Gamma(\frac{1}{2}+p-{\rm i}s)}{\Gamma(\frac{1}{2}+p+{\rm i}s)}\bigg{)}^{N({\tt d}\,|\,{\cal S})}\,\prod_{\bm{w}\atop{\tt d}{\rm\,-\,fixed}}\frac{\check{\mathfrak{C}}_{p,s}^{(+)}(\bm{w})}{\check{\mathfrak{C}}^{(-)}_{p,s}(\bm{w})}\,\Bigg{]}$ (4.5)
@@
up to corrections which vanish as $L\to\infty$. The product over $\bm{w}$ appearing in the r.h.s. goes over all the $N({\tt d}\,|\,\mathcal{S})$ (3.30) solution sets of the algebraic system (3.2) with ${\tt d}$ fixed. Also, recall that $2p=2{\cal S}+1-(n+2)$.
@@
In the work [45] a formula is presented for a product over $\bm{w}$ similar to the one appearing in the r.h.s. of (4.5) (see also Appendix B of [15]). It is valid for the case of generic $p$ and $n$ when the number of solution sets $\bm{w}$ of (3.2) is ${\rm par}_{2}({\tt d})$. Based on this, one can derive the result:
@@
$\prod_{\bm{w}\atop{\tt d}{\rm\,-\,fixed}}\frac{\check{\mathfrak{C}}_{p,s}^{(+)}(\bm{w})}{\check{\mathfrak{C}}^{(-)}_{p,s}(\bm{w})}$ $=$ $(-1)^{{\rm par}_{2}({\tt d}-2{\cal S}-1)}\prod_{a=0}^{{\tt d}-1}\bigg{(}\frac{\frac{1}{2}+a+p-{\rm i}s}{\frac{1}{2}+a+p+{\rm i}s}\bigg{)}^{N({\tt d}\,|\,{\cal S})-N_{a}^{+}({\tt d}\,|\,{\cal S})}$ (4.6)
$\times$ $\prod_{a=0}^{{\tt d}-1}\bigg{(}\frac{\frac{1}{2}+a-p-{\rm i}s}{\frac{1}{2}+a-p+{\rm i}s}\bigg{)}^{N({\tt d}\,|\,{\cal S})-N_{a}^{-}({\tt d}\,|\,{\cal S})}$
@@
$\sum_{{\tt d}=0}^{\infty}N_{a}^{\pm}({\tt d}\,|\,{\cal S})\,{\tt q}^{\tt d}=\bigg{(}\prod_{j=1}^{\infty}(1-{\tt q}^{j})^{-2}\bigg{)}\,\sum_{m=0}^{\infty}(-1)^{m}\,\big{(}1-{\tt q}^{(1\pm m)(2{\cal S}+1)}\big{)}\,{\tt q}^{ma+\frac{m(m+1)}{2}}\ .$ (4.7)
@@
$N_{a}^{+}({\tt d}\,|\,{\cal S})=N({\tt d}\,|\,{\cal S})-N_{-a-1}^{-}({\tt d}\,|\,{\cal S})\,.$ (4.8)
The scaling limit of the RG trajectory $L\mapsto|\Psi_{L}\rangle\in{\cal H}_{L|{\cal S}}^{({\rm cont})}$ labeled by real $s$, the integers ${\cal S}$, $S^{z}$, ${\tt d}$ and the solution set $\bm{w}$ yields
$\operatorname*{slim}_{L\to\infty}|\Psi_{L}\rangle=|\psi_{p,s}^{(S^{z})}(\bm{w})\rangle\ .$ (4.9)
@@
${\cal H}_{{\cal S},{\cal S}}^{({\rm cont})}=\int^{\oplus}_{\mathbb{R}}\mbox{d}s\ \bigoplus_{{\tt d}=0}^{\infty}{\cal H}^{({\rm cont},{\tt d})}_{p,s}\qquad\qquad\qquad\qquad\qquad\big{(}2p=2{\cal S}+1-n-2\big{)}\,.$ (4.11)
@@
$\dim\big{(}{\cal H}_{p,s}^{({\rm cont},{\tt d})}\big{)}=N({\tt d}\,|\,\mathcal{S})\ .$ (4.12)
For the low energy states where the value of $\Im m\big{(}b(L)\big{)}$ is nonvanishing in the limit $L\to\infty$ so that they do not belong to ${\cal H}_{L|{\cal S}}^{({\rm cont})}$, one may repeat the similar analysis that was performed in ref.[15] in the context of the staggered six-vertex model with quasi-periodic BCs (see Appendix B therein). Let’s denote by ${\cal H}_{L|{\cal S},S^{z}}^{({\rm disc})}$ the set of such states $|\Psi_{L}\rangle$ with given quantum numbers ${\cal S}$ and $S^{z}$. The quantization condition (3.40) implies that the set $\bm{w}$ and $s=\operatorname*{slim}_{L\to\infty}b(L)$ labelling the RG trajectory $\{|\Psi_{L}\rangle\}$ must be such that
$\mbox{e}^{-\frac{{\rm i}}{2}\delta(\bm{w},p,s)}=0\qquad{\rm if}\qquad\Im m(s)>0\,,\qquad\qquad\qquad\mbox{e}^{+\frac{{\rm i}}{2}\delta(\bm{w},p,s)}=0\qquad{\rm if}\qquad\Im m(s)<0\,.$ (4.13)
@@
$\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w},p,s)}=\mbox{e}^{-\frac{{\rm i}}{2}\delta(-\bm{w},p,-s)}\,.$ (4.15)
Here $-\bm{w}$ denotes the set $\{-w_{a}\}_{a=1}^{\tt d}$, where if $\bm{w}$ obeys the algebraic system (3.2) then $-\bm{w}$ obeys the same set of equations with $s\mapsto-s$. This allows one to focus on the case with $0<\Im m(s)\leq\frac{n}{2}$, while results for $-\frac{n}{2}<\Im m(s)<0$ follow by simply flipping the sign $s\mapsto-s$.
@@
$\prod_{\bm{w}\atop{\tt d}{\rm\,-\,fixed}}\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w},p,s)}=\bigg{(}2^{(2n+4)\frac{{\rm i}s}{n}}\ \frac{\Gamma(\frac{1}{2}+p-{\rm i}s)}{\Gamma(\frac{1}{2}+p+{\rm i}s)}\bigg{)}^{N({\tt d}\,|\,{\cal S})}\,\prod_{\bm{w}\atop{\tt d}{\rm\,-\,fixed}}\frac{\check{\mathfrak{C}}_{p,s}^{(+)}(\bm{w})}{\check{\mathfrak{C}}^{(-)}_{p,s}(\bm{w})}$ (4.16)
with the last term in the r.h.s. being given by the product (4.6). It follows from the definitions (3.41), (3.33) and (3.34). Also, we’ll need the following assumptions on the positions of the poles and zeroes of the function $\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w},p,s)}$, which were verified numerically for small ${\tt d}\leq 3$:
@@
    The points where $\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w},p,s)}$ is singular do not coincide with the location of any zero of $\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w}^{\prime},p,s)}$ with $\bm{w}^{\prime}$ being some other solution set of (3.2).
@@
    All singularities of $\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w},p,s)}$ in the complex $s$ plane are simple poles. Notice that, in view of eq. (4.15), this implies that all of its zeroes are simple as well.
From assumption (i), any pole or zero of $\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w},p,s)}$ must appear as a pole/zero in the r.h.s. of (4.16). This way, one finds that the values of $s$ for which the first condition in (4.13) is obeyed are $s=\pm s_{a}$, with
@@
where the bound on $a+{\cal S}$ comes from (4.14). Moreover, due to (ii) one can determine the number of solution sets $\bm{w}$ with (4.13) being satisfied at $s=s_{a}$ by reading off the multiplicity of that pole/zero from eqs. (4.16) and (4.6). This would coincide with the dimension of the linear space ${\cal H}_{p,s}^{({\rm disc},{\tt d})}$, which is the span of all states of the form $|\psi_{p,s}^{(S^{z})}(\bm{w})\rangle$ having fixed ${\cal S}$, $S^{z}$, ${\tt d}$ and $s$ with $\Im m(s)\neq 0$. One finds the number of such $\bm{w}$ to be $N_{a}^{+}({\tt d})$.
@@
${\cal H}_{{\cal S},{\cal S}}^{({\rm disc})}=\bigoplus_{s\in\Sigma^{+}\cup\Sigma^{-}}\ \bigoplus_{{\tt d}=0}^{\infty}{\cal H}_{p,s}^{({\rm disc},{\tt d})}\ .$ (4.18)
Here $\Sigma(\pm)$ denote the finite sets of pure imaginary numbers:
@@
which incorporate the bound on the imaginary part of $s$ (4.14). Each component ${\cal H}_{p,s}^{({\tt d})}$ is finite dimensional and
$\dim\big{(}{\cal H}_{p,s}^{({\rm disc},{\tt d})}\big{)}=N^{+}_{a}({\tt d}\,|\,{\cal S})\qquad\qquad{\rm with}\qquad\qquad a=\tfrac{n}{2}-{\cal S}\pm{\rm i}s\in\mathbb{Z}$ (4.20)
@@
The following comment is in order here. For the case $a<0$, the integers $N_{a}^{+}({\tt d}\,|\,{\cal S})$ (4.7) are all zero for ${\tt d}=0$:

### v3

@@ 4.1 Continuous and discrete spectrum
A key result of this paper is the conjecture, above, which was motivated by our numerical work. It describes a certain one-to-one relation between $b(L)$ and $b_{*}(L)$. The former is obtained via the Bethe roots corresponding to a state $|\Psi_{L}\rangle$ in a multiplet ${\cal M}_{\cal S}$, labeled by $2p=2{\cal S}+1-(n+2)$, the non-negative integer d and one of the $N(\texttt{d}\,|\,{\cal S})={\rm par}_{2}(\texttt{d})-{\rm par}_{2}(\texttt{d}-2{\cal S}-1)$ solution sets $\boldsymbol{w}=\{w_{a}\}_{a=1}^{\texttt{d}}$ of the algebraic system (3.2). Since it requires the construction of an RG trajectory, computing $b(L)$ for $L\gg 1$ can be cumbersome and time-consuming to carry out. The notation $b_{*}(L)$ stands for a solution of the quantization condition (3.41) treated as an equation for $b(L)$ with the correction terms ignored. Accepting the conjecture to be true, one can determine the spectrum of $b(L)$ for the low energy states at large $L$ via a study of (3.41). This is extremely powerful, since much less computing resources are needed to numerically analyze the quantization condition and it may be studied analytically as well. The results allow one to characterize the spectrum of conformal dimensions together with the space of low energy states in the scaling limit.
Consider the quantization condition (3.41) and suppose that $\delta(\boldsymbol{w},p,s)|_{s\mapsto b(L)}\ll\log(L)$. Then the first term dominates and one can develop an asymptotic expansion for $b(L)$ in $1/\log(L)$. The leading and subleading asymptotic behaviour reads as
@@
Each low energy Bethe state in ${\cal H}_{L|{\cal S},S^{z}}^{({\rm cont})}$ is uniquely specified by the non-negative integer d, a solution set $\boldsymbol{w}=\{w_{a}\}_{a=1}^{\texttt{d}}$ of the algebraic system (3.2) with $s\mapsto b(L)$, as well as the even or odd integer $\mathfrak{m}$ that enters into the asymptotics (4.1). For $L\gg 1$ the value of $b_{\mathfrak{m}}(L)$ becomes densely distributed in the segment $(-b_{\mathfrak{m}_{\rm max}}(L),+b_{\mathfrak{m}_{\rm max}}(L))$. The density of states is obtained from the quantization condition (3.41) written in logarithmic form:
$\big{[}4s\log(L/L_{0})+\delta(\boldsymbol{w},p,s)\big{]}|_{s\mapsto b_{\mathfrak{m}}(L)}=2\pi\mathfrak{m}+O(L^{-\epsilon})\ .$ (4.4)
@@
Here the branch of the logarithm needed to define $\delta$ from the relation (3.41b) is taken such that the phase shift is a continuous function of $s$ in the strip $|\Im m(s)|<\varepsilon$ for some $\varepsilon>0$ (it is being assumed that $\mbox{e}^{\frac{{\rm i}}{2}\delta}$ contains no zeroes or poles for real $s$). The term in the square brackets in the left hand side of (4.4) is a monotonic function of $s$ for $L$ sufficiently large. This way one concludes that the number of states in ${\cal H}_{L|{\cal S},S^{z}}^{({\rm cont})}$ with fixed d such that $\Re e\big{(}b(L)\big{)}$ lies in the interval $(s,s+\Delta s)\in(-b_{\mathfrak{m}_{\rm max}}(L),+b_{\mathfrak{m}_{\rm max}}(L))$ is given by $\rho^{(\texttt{d})}_{p}(s)\,\Delta s$ with [^6] This line of arguments is analogous to the standard derivation in the root density approach. One introduces a monotonic increasing counting function which evaluates to (half-)integers at the Bethe roots similar as the l.h.s of (4.4) evaluates to odd/even integers $\mathfrak{m}$ multiplied by $2\pi$ when $s$ is swapped for $b_{\mathfrak{m}}(L)$ and $L\gg 1$. Differentiating the counting function in the root density approach yields the root density, while we obtain the density of states (4.5) by differentiating (4.4) and dividing by $4\pi$.
@@
$\rho^{(\texttt{d})}_{p}(s)=\frac{1}{\pi}\,N(\texttt{d}\,|\,{\cal S})\,\log\big{(}2^{\frac{n+2}{n}}\,L/L_{0}\big{)}+\frac{1}{2\pi{\rm i}}\ \partial_{s}\log\Bigg{[}\bigg{(}\frac{\Gamma(\frac{1}{2}+p-{\rm i}s)}{\Gamma(\frac{1}{2}+p+{\rm i}s)}\bigg{)}^{N(\texttt{d}\,|\,{\cal S})}\,\prod_{\boldsymbol{w}\atop\texttt{d}{\rm\,-\,fixed}}\frac{\check{\mathfrak{C}}_{p,s}^{(+)}(\boldsymbol{w})}{\check{\mathfrak{C}}^{(-)}_{p,s}(\boldsymbol{w})}\,\Bigg{]}$ (4.5)
@@
up to corrections which vanish as $L\to\infty$. The product over $\boldsymbol{w}$ appearing in the r.h.s. goes over all the $N(\texttt{d}\,|\,\mathcal{S})$ (3.31) solution sets of the algebraic system (3.2) with d fixed. Also, recall that $2p=2{\cal S}+1-(n+2)$.
@@
In the work [45] a formula is presented for a product over $\boldsymbol{w}$ similar to the one appearing in the r.h.s. of (4.5) (see also Appendix B of [15]). It is valid for the case of generic $p$ and $n$ when the number of solution sets $\boldsymbol{w}$ of (3.2) is ${\rm par}_{2}(\texttt{d})$. Based on this, one can derive the result:
@@
$\prod_{\boldsymbol{w}\atop\texttt{d}{\rm\,-\,fixed}}\frac{\check{\mathfrak{C}}_{p,s}^{(+)}(\boldsymbol{w})}{\check{\mathfrak{C}}^{(-)}_{p,s}(\boldsymbol{w})}$ $=$ $(-1)^{{\rm par}_{2}(\texttt{d}-2{\cal S}-1)}\prod_{a=0}^{\texttt{d}-1}\bigg{(}\frac{\frac{1}{2}+a+p-{\rm i}s}{\frac{1}{2}+a+p+{\rm i}s}\bigg{)}^{N(\texttt{d}\,|\,{\cal S})-N_{a}^{+}(\texttt{d}\,|\,{\cal S})}$ (4.6)
$\times$ $\prod_{a=0}^{\texttt{d}-1}\bigg{(}\frac{\frac{1}{2}+a-p-{\rm i}s}{\frac{1}{2}+a-p+{\rm i}s}\bigg{)}^{N(\texttt{d}\,|\,{\cal S})-N_{a}^{-}(\texttt{d}\,|\,{\cal S})}$
@@
$\sum_{\texttt{d}=0}^{\infty}N_{a}^{\pm}(\texttt{d}\,|\,{\cal S})\,\texttt{q}^{\texttt{d}}=\bigg{(}\prod_{j=1}^{\infty}(1-\texttt{q}^{j})^{-2}\bigg{)}\,\sum_{m=0}^{\infty}(-1)^{m}\,\big{(}1-\texttt{q}^{(1\pm m)(2{\cal S}+1)}\big{)}\,\texttt{q}^{ma+\frac{m(m+1)}{2}}\ .$ (4.7)
@@
$N_{a}^{+}(\texttt{d}\,|\,{\cal S})=N(\texttt{d}\,|\,{\cal S})-N_{-a-1}^{-}(\texttt{d}\,|\,{\cal S})\,.$ (4.8)
The scaling limit of the RG trajectory $L\mapsto|\Psi_{L}\rangle\in{\cal H}_{L|{\cal S}}^{({\rm cont})}$ labeled by real $s$, the integers ${\cal S}$, $S^{z}$, d and the solution set $\boldsymbol{w}$ yields
$\operatorname*{slim}_{L\to\infty}|\Psi_{L}\rangle=|\psi_{p,s}^{(S^{z})}(\boldsymbol{w})\rangle\ .$ (4.9)
@@
${\cal H}_{{\cal S},{\cal S}}^{({\rm cont})}=\int^{\oplus}_{\mathbb{R}}\mbox{d}s\ \bigoplus_{\texttt{d}=0}^{\infty}{\cal H}^{({\rm cont},\texttt{d})}_{p,s}\qquad\qquad\qquad\qquad\qquad\big{(}2p=2{\cal S}+1-n-2\big{)}\,.$ (4.11)
@@
$\dim\big{(}{\cal H}_{p,s}^{({\rm cont},\texttt{d})}\big{)}=N(\texttt{d}\,|\,\mathcal{S})\ .$ (4.12)
For the low energy states where the value of $\Im m\big{(}b(L)\big{)}$ is nonvanishing in the limit $L\to\infty$ so that they do not belong to ${\cal H}_{L|{\cal S}}^{({\rm cont})}$, one may repeat the similar analysis that was performed in ref.[15] in the context of the staggered six-vertex model with quasi-periodic BCs (see Appendix B therein). Let’s denote by ${\cal H}_{L|{\cal S},S^{z}}^{({\rm disc})}$ the set of such states $|\Psi_{L}\rangle$ with given quantum numbers ${\cal S}$ and $S^{z}$. The quantization condition (3.41) implies that the set $\boldsymbol{w}$ and $s=\operatorname*{slim}_{L\to\infty}b(L)$ labelling the RG trajectory $\{|\Psi_{L}\rangle\}$ must be such that
$\mbox{e}^{-\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}=0\qquad{\rm if}\qquad\Im m(s)>0\,,\qquad\qquad\qquad\mbox{e}^{+\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}=0\qquad{\rm if}\qquad\Im m(s)<0\,.$ (4.13)
@@
$\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}=\mbox{e}^{-\frac{{\rm i}}{2}\delta(-\boldsymbol{w},p,-s)}\,.$ (4.15)
Here $-\boldsymbol{w}$ denotes the set $\{-w_{a}\}_{a=1}^{\texttt{d}}$, where if $\boldsymbol{w}$ obeys the algebraic system (3.2) then $-\boldsymbol{w}$ obeys the same set of equations with $s\mapsto-s$. This allows one to focus on the case with $0<\Im m(s)\leq\frac{n}{2}$, while results for $-\frac{n}{2}<\Im m(s)<0$ follow by simply flipping the sign $s\mapsto-s$.
@@
$\prod_{\boldsymbol{w}\atop\texttt{d}{\rm\,-\,fixed}}\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}=\bigg{(}2^{(2n+4)\frac{{\rm i}s}{n}}\ \frac{\Gamma(\frac{1}{2}+p-{\rm i}s)}{\Gamma(\frac{1}{2}+p+{\rm i}s)}\bigg{)}^{N(\texttt{d}\,|\,{\cal S})}\,\prod_{\boldsymbol{w}\atop\texttt{d}{\rm\,-\,fixed}}\frac{\check{\mathfrak{C}}_{p,s}^{(+)}(\boldsymbol{w})}{\check{\mathfrak{C}}^{(-)}_{p,s}(\boldsymbol{w})}$ (4.16)
with the last term in the r.h.s. being given by the product (4.6). It follows from the definitions (3.41b), (3.34) and (3.35). Also, we’ll need the following assumptions on the positions of the poles and zeroes of the function $\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}$, which were verified numerically for small $\texttt{d}\leq 3$:
@@
    The points where $\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}$ is singular do not coincide with the location of any zero of $\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w}^{\prime},p,s)}$ with $\boldsymbol{w}^{\prime}$ being some other solution set of (3.2).
@@
    All singularities of $\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}$ in the complex $s$ plane are simple poles. Notice that, in view of eq. (4.15), this implies that all of its zeroes are simple as well.
From assumption (i), any pole or zero of $\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}$ must appear as a pole/zero in the r.h.s. of (4.16). This way, one finds that the values of $s$ for which the first condition in (4.13) is obeyed are $s=\pm s_{a}$, with
@@
where the bound on $a+{\cal S}$ comes from (4.14). Moreover, due to (ii) one can determine the number of solution sets $\boldsymbol{w}$ with (4.13) being satisfied at $s=s_{a}$ by reading off the multiplicity of that pole/zero from eqs. (4.16) and (4.6). This would coincide with the dimension of the linear space ${\cal H}_{p,s}^{({\rm disc},\texttt{d})}$, which is the span of all states of the form $|\psi_{p,s}^{(S^{z})}(\boldsymbol{w})\rangle$ having fixed ${\cal S}$, $S^{z}$, d and $s$ with $\Im m(s)\neq 0$. One finds the number of such $\boldsymbol{w}$ to be $N_{a}^{+}(\texttt{d})$.
@@
${\cal H}_{{\cal S},{\cal S}}^{({\rm disc})}=\bigoplus_{s\in\Sigma^{+}\cup\Sigma^{-}}\ \bigoplus_{\texttt{d}=0}^{\infty}{\cal H}_{p,s}^{({\rm disc},\texttt{d})}\ .$ (4.18)
Here $\Sigma^{\pm}$ denote the finite sets of pure imaginary numbers:
@@
which incorporate the bound on the imaginary part of $s$ (4.14). Each component ${\cal H}_{p,s}^{(\texttt{d})}$ is finite dimensional and
$\dim\big{(}{\cal H}_{p,s}^{({\rm disc},\texttt{d})}\big{)}=N^{+}_{a}(\texttt{d}\,|\,{\cal S})\qquad\qquad{\rm with}\qquad\qquad a=\tfrac{n}{2}-{\cal S}\pm{\rm i}s\in\mathbb{Z}$ (4.20)
@@
The following comment is in order here. For the case $a<0$, the integers $N_{a}^{+}(\texttt{d}\,|\,{\cal S})$ (4.7) are all zero for $\texttt{d}=0$: